OpenStudy (jaylelile):

8.Sebastian writes the recursive formula f(x+1) = 4f(x) to represent a geometric sequence whose second term is 12. Which explicit formula can be used to model the same sequence? f(x) = 12(4)x f(x) = 3(4)x − 1 f(x) = 4(12)x f(x) = 4(3)x − 1

OpenStudy (misssunshinexxoxo):

Which one do you believe it is?

OpenStudy (jaylelile):

I come up with different answers every time

OpenStudy (anonymous):

If f(2)=12 then f(2+1)=4(12) Now go to the general formula for a geometric sequence. \[a _{n}=a _{1}*r ^{(n-1)}\] Since f(3)=48 and f(2)=12 or you can say \[a _{3}=48\] and \[a _{2}=12\] Since \[r=\frac{ a _{n +1} }{ a _{n} }\]then r= 48/12 or 4 Now find \[a _{1}\] from the formula for n=3 \[48=a _{1}*4^{(3-1)}\] So \[a _{1}=3\] So it looks like \[f(x)=3*4^{x-1}\]